Circular No. 8234

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN M33 G. J. Schwarz, Steward Observatory; R. M. Wagner, LBT Observa-

tory; S. Starrfield, Arizona State University; and P. Szkody, University of Washington, report that spectroscopy of the possible nova reported on IAUC 8195, with the 6.5-m MMT (range 400–900 nm; resolution 0.8 nm) on Sept. 13.44 UT, confirms the object to be a classical nova in its early decline phase. The spectra exhibit strong emission lines of the Balmer series of hydrogen; numerous Fe II blends (multiplets 37, 38, 42, 48, 49, and 74); O I at 777.4 and 844.6 nm; and the Ca II infrared triplet at 854.2, 866.2, and 849.8 nm that is typical of ‘Fe II’-class novae after maximum light. The equivalent width of Hα emission is 40 nm, and it exhibits a pronounced castellated structure with a FWZI of almost 5400 km/s.

SUPERNOVAE 2003jd, 2003je, AND 2003jf A. V. Filippenko, R. T. Foley, and B. Swift, University of California,

report that inspection of CCD spectra (range 320–1000 nm), obtained on Oct. 28 UT with the Shane 3-m telescope at Lick Observatory, reveals that SN 2003jd (IAUC 8232) is probably a peculiar type-Ic supernova, with some similarities to SNe 1998bw and 2002ap (Foley et al. 2003, PASP 115, 1220). The expansion velocity, measured from the bottom of the putative Ca II near-infrared triplet absorption line, is ∼ 40000 km/s. The continuum is quite blue, perhaps more so than that of other 1998bw-like objects at a similar phase. It is possible that the object is a 1991T-like type-Ia supernova, although the peculiar type-Ic hypothesis is currently favored. A spectrum of SN 2003je (IAUC 8232) shows it to be of type II (perhaps type IIn), with low-contrast Hα emission superposed on a nearly featureless blue continuum. The corresponding absorption component of Hα is weak or absent. A noisy spectrum of SN 2003jf (IAUC 8233) suggests that it is a young type-II supernova with a featureless blue continuum.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003jd, obtained by P. Berlind on Oct. 28.32 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a peculiar type-Ic supernova. The spectrum shows some similarities with SN 1998bw near maximum (Patat et al. 2001, Ap.J. 555, 900), but it is not identical. Further observations are encouraged.

2003 October 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green